# Transic: Sim-to-Real Policy Transfer by Learning from Online Correction

### v1

@@ 3.4 Implementation Details
We use Isaac Gym [10] as the simulation backend. Proximal policy optimization (PPO [84]) is used to train teacher policies from scratch. We design task-specific reward functions and curricula when necessary to facilitate RL training. We apply exhaustive domain randomization during teacher policy training and proper data augmentation during student policy distillation. Student policies are parameterized as Gaussian Mixture Models (GMMs [68]). We have also experimented with other state-of-the-art policy models, such as Diffusion Policy [85], but did not observe better performances. See the Appendix Sec. A for more details about the simulation training phase and additional comparisons. During the human-in-the-loop data collection phase, we use a 3Dconnexion SpaceMouse as the teleoperation interface. Residual policies use state-of-the-art point cloud encoders, such as PointNet [86] and Perceiver [87, 88], and GMM as the action head. We follow the best practices to train residual policies, including using learning rate warm-up and cosine annealing [89]. More training hyperparameters are provided in the Appendix Sec. B.4.
@@ 4 Experiments
(a) Stabilize
(b) Reach and Grasp
(c) Insert
(d) Screw
Figure 3: Four tasks benchmarked in this work. They are fundamental skills required to assemble a square table from FurnitureBench [90]. We randomize objects’ initial poses during evaluation. a) The robot pushes the square tabletop to the right corner of the wall such that it remains stable in following assembly steps. b) The robot reaches and grasps the table leg. It needs to properly adjust the end effector’s orientation to avoid infeasible grasping poses. c) The robot inserts the pre-grasped table leg to the far right assembly hole of the tabletop. d) The robot’s end-effector is initialized close to an inserted table leg and it screws the table leg clockwise into the tabletop.
We seek to answer the following research questions with our experiments:
@@ 1. $\mathbf{\mathcal{Q}1}$:
    Does Transic lead to better transfer performance compared to traditional sim-to-real methods (Sec. 4.2)?
@@ 2. $\mathbf{\mathcal{Q}2}$:
    Can Transic better integrate human correction into the policy learned in simulation than existing interactive imitation learning (IL) approaches (Sec. 4.2)?
@@ 3. $\mathbf{\mathcal{Q}3}$:
    Does Transic require less real-world data to achieve good performance compared to algorithms that only learn from real-robot trajectories (Sec. 5)?
@@ 4. $\mathbf{\mathcal{Q}4}$:
    How effective can Transic address different types of sim-to-real gaps (Sec. 4.3)?
  5. $\mathbf{\mathcal{Q}5}$:
    How does Transic scale with human effort (Sec. 4.4)?
  6. $\mathbf{\mathcal{Q}6}$:
@@
    Does Transic exhibit intriguing properties, such as generalization to unseen objects, effective gating, policy robustness, consistency in learned visual features, ability to solve long-horizon manipulation tasks, and other emergent behaviors (Sec. 4.5)?
4.1 Experiment Settings
Tasks
We consider complex contact-rich manipulation tasks that require high precision in FurnitureBench [90]. These tasks are challenging and ideal for testing sim-to-real transfer, since perception, embodiment, controller, and dynamics gaps all need to be addressed to accomplish the tasks successfully. Specifically, we divide the assembly of a square table into four independent tasks (Fig. 3): Stabilize, Reach and Grasp, Insert, and Screw. We collect 20, 100, 90, and 17 real-robot trajectories with human correction, respectively. To further test the generalization to unseen objects from a new category, we experiment with a lamp (Fig. 6(b)). All experiments are conducted on a tabletop setting with a mounted Franka Emika 3 robot. See the Appendix Sec. B.1 for the detailed system setup.
@@ 4.1 Experiment Settings
Figure 4: Average success rates over four benchmarked tasks. Transic significantly outperforms three baseline groups. Results are success rates averaged over four tasks.
Baselines and Evaluation Protocol
We compare with the following three groups of baselines. 1) Traditional sim-to-real methods: This group includes direct deployment of simulation policy trained with domain randomization and data augmentation [53], denoted as “DR. & Data Aug.”. It also covers the real-world fine-tuning paradigm, where simulation policies are further fine-tuned with real-robot data through BC (denoted as “BC Fine-Tune”) and the state-of-the-art offline RL method (implicit Q-learning [69], denoted as “IQL Fine-Tune”). To estimate the performance lower bound, we also include the baseline without any data augmentation or real-world fine-tuning, denoted as “Direct Transfer”. 2) Interactive IL: This group represents the state-of-the-art interactive imitation learning methods, including HG-Dagger [66] and IWR [67]. 3) Learning from real-robot data only: This group includes BC [72], BC-RNN [68], and IQL [69]. They are trained on real-robot demonstrations only. We follow Liu et al. [70] to label reward for IQL. All evaluations consist of 20 trials starting with different objects and robot poses. We make our best efforts to ensure the same initial settings when evaluating different methods. See Appendix Sec. C for the detailed evaluation protocol.
Table 1: Success rates per tasks. Transic outperforms all baseline methods on all four tasks.
Tasks Transic
Direct
Transfer
DR. & Data
Aug. [53]
BC
Fine-Tune
IQL
Fine-Tune
HG-Dagger [66] IWR [67] BC [72] BC-RNN [68] IQL [69]
Stabilize 100% 10% 35% 55% 0% 65% 65% 40% 40% 5%
Reach and Grasp 95% 35% 60% 35% 0% 30% 40% 25% 0% 5%
Insert 45% 0% 15% 15% 25% 35% 40% 10% 5% 0%
Screw 85% 0% 35% 50% 65% 40% 40% 15% 25% 0%
4.2 Quantitative Comparison on Four Assembly Tasks
As shown in Fig. 4 and Table 1, Transic achieves the best performance on average and on all four tasks with significant margins. We now compare them in detail and discuss the main findings.
Transic is effective for sim-to-real transfer ($\mathbf{\mathcal{Q}1}$). It successfully achieves a high average success rate of 81% across four tasks on real robots. For the task Stabilize, it can achieve a 100% success rate while the best baseline method (BC Fine-Tune) only succeeds half the time. For challenging tasks that require precise and contact-rich manipulation, such as Insert and Screw, Transic results in promising outcomes (45% and 85% success rates, respectively) while the Direct Transfer baseline never succeeds.
@@ 4.2 Quantitative Comparison on Four Assembly Tasks
What are the reasons for successful transfer? We observe that adding real-world human correction data does not guarantee improvement. For example, among traditional sim-to-real methods, the best baseline BC Fine-Tune outperforms DR. & Data Aug. by 7% but IQL Fine-Tune leads to worse performance. In contrast, Transic effectively uses human correction data, which boosts average performance by 124%. Presumably, BC Fine-Tune’s marginal improvement is due to the domain difference between simulation and reality. This significant distinction cannot be easily bridged through naïve fine-tuning. Overall, Transic not only achieves the best transfer performance, but also improves simulation policies the most among various sim-to-real approaches.
@@
Transic can better incorporate human correction into the original learned policy ($\mathbf{\mathcal{Q}2}$). Due to the differences between human and robot behaviors [68], using real-world data to directly fine-tune a policy that has been largely trained on machine-generated trajectories can lead to undesired performance. Specifically, Transic outperforms interactive IL methods including HG-Dagger [66] and IWR [67] by 75% on average. While both of them weigh the intervention data higher during training (Sec. 2.2), we find that they tend to erase the original policy and lead to catastrophic forgetting. Therefore, in state space where no human intervention exists, they behave suboptimally and hence suffer from out-of-distribution issues due to compounding error. In contrast, by incorporating human correction with a separate residual policy and integrating both base and residual policies through gating, Transic combines the best properties of both policies during deployment. It relies on the simulation policy for robust execution most of the time; when the base policy is likely to fail, it automatically applies the residual policy to prevent failures and correct mistakes.
@@
Transic requires significantly less real-world data ($\mathbf{\mathcal{Q}3}$). It only requires dozens of real-robot trajectories to achieve superior performance. However, methods such as BC-RNN and IQL trained on such a limited number of data suffer from overfitting and model collapse. Transic achieves 3.6$\times$ better performance than them. In fact, as task complexity increases, we may need exponentially more real-world data to train these models, which solely rely on real-world demonstration data, according to the previous literature [68]. This result highlights the importance of training in simulation first, and then leveraging sim-to-real transfer for robot learning practitioners.
@@
Summary we show that in sim-to-real transfer, a good base policy learned from the simulation can be combined with limited real-world data to achieve success ($\mathbf{\mathcal{Q}3}$). However, effectively utilizing human correction data to address the sim-to-real gap is challenging ($\mathbf{\mathcal{Q}1}$), especially when we want to prevent catastrophic forgetting of the base policy ($\mathbf{\mathcal{Q}2}$).
Figure 6: a) Scalability with human correction data. Numbers are success rates averaged over four tasks with different amount of human correction data. Per-task results are shown in Table A.XI. b) Generalization to unseen objects from a new category. Success rates are averaged over tasks Reach and Grasp and Screw.
@@
4.3 Effectiveness in Addressing Different Sim-to-Real Gaps ($\mathbf{\mathcal{Q}4}$)
@@
While Transic is a holistic approach to address multiple sim-to-real gaps simultaneously, we shed light on its ability to close each individual gap. To do so, we create five different simulation-reality pairs. For each of them, we intentionally create large gaps between the simulation and the real world. These gaps are applied to the real-world setting and they include perception error, underactuated controller, embodiment mismatch, dynamics difference, and object asset mismatch. Note that these are artificial settings for a controlled study. See the Appendix Sec. C.2 for detailed setups.
@@
As shown in Fig. 5, Transic achieves an average success rate of 77% across five different simulation-reality pairs with deliberately exacerbated sim-to-real gaps. This indicates its remarkable ability to close these individual gaps. In contrast, the best baseline method, IWR, only achieves an average success rate of 18%. We attribute this effectiveness in addressing different sim-to-real gaps to the residual policy design. Zeng et al. [83] echos our finding that residual learning is an effective tool to compensate for domain factors that cannot be explicitly modeled. Furthermore, training with data specifically collected from a particular setting generally increases Transic’s performance. However, this is not the case for IWR, where fine-tuning on new data can even lead to worse performance. These results show that Transic is better not only in addressing multiple sim-to-real gaps as a whole, but also in handling individual types of gaps of very different nature.
4.4 Scalability with Human Effort ($\mathbf{\mathcal{Q}5}$)
@@
Scaling with human effort is a desired property for human-in-the-loop robot learning methods [70]. We show that Transic has better human data scalability than the best baseline IWR in Fig. 6(a) and Table A.XI. If we increase the size of the correction dataset from 25% to 75% of the full dataset size, Transic achieves a relative improvement of 42% in the average success rate. In contrast, IWR only achieves 23% relative improvement. Additionally, for tasks other than Insert, IWR performance plateaus at an early stage and even starts to decrease as more human data becomes available. We hypothesize that IWR suffers from catastrophic forgetting and struggles to properly model the behavioral modes of humans and trained robots. On the other hand, Transic bypasses these issues by learning gated residual policies only from human correction.
4.5 Intriguing Properties and Emergent Behaviors ($\mathbf{\mathcal{Q}6}$)
Finally, we examine further Transic and discuss several emergent capabilities.
Generalization to Unseen Objects
We show that a robot trained with Transic can zero-shot generalize to new objects from a novel category. As shown in Fig. 6(b), Transic can achieve an average success rate of 75% when zero-shot evaluated on assembling a lamp. However, IWR can only succeed once every three attempts. This evidence suggests that Transic is not overfitted to a particular object, instead, it has learned reusable skills for category-level object generalization.
Effects of Different Gating Mechanisms
We introduce the learned gated residual policy in Sec. 3.3 where the gating mechanism controls when to apply residual actions. To assess the quality of learned gating, we compare its performance with an actual human operator performing gating. Results are shown in Table 2. It is evident that the learned gating mechanism only incurs negligible performance drops compared to human gating. This suggests that Transic can reliably operate in a fully autonomous setting once the gating mechanism is learned.
Table 2: Results of ablation studies. We study the effects of different gating mechanisms (learned gating vs human gating), policy robustness against reduced cameras and suboptimal correction data, and the importance of visual encoder regularization.
@@ 4.5 Intriguing Properties and Emergent Behaviors ($\mathbf{\mathcal{Q}6}$)
Policy Robustness
We investigate the policy robustness against 1) point cloud observations with inferior quality by removing two cameras, and 2) suboptimal correction data with noise injection. See Appendix Sec. C.4 for detailed experiment setups. Results are shown in Table 2. We highlight that Transic is robust to partial point cloud inputs caused by the reduced number of cameras. We attribute this to the heavy point cloud downsampling employed during training. Fishman et al. [91] echos our finding that policies trained with downsampled synthetic point cloud inputs can generalize to partial point cloud observations obtained in the real world without the need for shape completion. Meanwhile, when the correction data used to learn residual policies are suboptimal, Transic only shows a relative decrease of 6% in the average success rate. We attribute this to the advantage of our integrated deployment — when the residual policy behaves suboptimally, the base policy could still compensate for the error in subsequent steps.
Importance of Point Cloud Encoder Regularization
To learn consistent visual features between the simulation and reality, we propose to regularize the point cloud encoder during the distillation stage as in Eq. 1. As shown in Table 2, the performance significantly decreases without such regularization, especially for tasks that require fine-grained visual features. Without it, simulation policies would overfit to synthetic point cloud observations and hence are not ideal for sim-to-real transfer.
Qualitative Analysis and Emergent Behaviors
We first examine the distribution of the collected human correction dataset. During the human-in-the-loop data collection, the probability of intervening and correcting is reasonably low ($P_{\text{correction}}\approx 0.20$). This is consistent with our intuition that, with a good base policy, interventions are not necessary for most of the time. However, they become critical when the robot tends to behave abnormally due to unaddressed sim-to-real gaps. Moreover, as highlighted in Fig. A.8, interventions happen at different times across tasks. This fact renders heuristics-based methods [92] for deciding when to intervene difficult, and further necessitates our learned residual policy.
Surprisingly, Transic shows several representative behaviors that resemble humans. For instance, they include error recovery, unsticking, safety-aware actions, and failure prevention as shown in Fig. 7.
Solving Long-Horizon Manipulation Tasks
Finally, we demonstrate that successful sim-to-real transfer of individual skills can be effectively chained together to enable long-horizon contact-rich manipulation (Fig. 8). See videos on transic-robot.github.io for a robot assembling a square table and a lamp using Transic.
(a) Error Recovery
(b) Unsticking
(c) Safety-aware Actions
(d) Failure Prevention
Figure 7: Emergent behaviors learned by Transic. a) Error recovery. Left: The robot tries to insert the table leg but the direction is wrong; Right: Transic raises the end effector and moves to the correct insertion position. b) Unsticking. Left: The robot hovers for a while and never reaches the light bulb; Right: Transic helps the robot get unstuck and move to the bulb. c) Safety-aware actions. Left: When pushing the tabletop, the gripper is too low and bends. This might damage the robot; Right: Transic compensates for the command that causes the end effector to move too low. d) Failure prevention. Left: The light bulb will fall and break after gripper opening; Right: Transic adjusts the bulb to a stable pose to prevent failure.
(a) Assemble a lamp.
(b) Assemble a square table.
Figure 8: We demonstrate that individual skills learned with Transic can be effectively chained together to enable long-horizon contact-rich manipulation. a) The robot assembles a lamp (160 seconds in 1$\times$ speed). b) The robot assembles a square table from FurnitureBench [90] (550 seconds in 1$\times$ speed). Videos are available on transic-robot.github.io.
@@ 5 Related Work
Although these gaps may not be fully bridged, traditional methods to address them include system identification [51, 19, 52, 31], domain randomization [53, 14, 54, 55], real-world adaptation [56], and simulator augmentation [58, 59, 60]. However, system identification is mostly engineered on a case-by-case basis. Domain randomization suffers from the inability to identify and randomize all physical parameters. Methods with real-world adaptation, usually through meta-learning [114], incur potential safety concerns during the adaptation phase. Most of these approaches also rely on explicit and domain-specific knowledge about tasks and the simulator a priori. For instance, to perform system identification for closing the embodiment gap for a quadruped, Tan et al. [19] disassembles the physical robot and carefully calibrates parameters including size, mass, and inertia. Kim et al. [33] reports that collaborative robots, such as the commonly used Franka Emika robot, have intricate joint friction that is hard to identify and randomized in typical physics simulators. To make a simulator more akin to the real world, Chebotar et al. [40] deploys trained virtual robots multiple times to refine the distributions of simulation parameters. This procedure not only introduces a significant real-world sampling effort, but also incurs potential safety concerns due to deploying suboptimal policies. In contrast, our method leverages human intervention data to implicitly overcome the transferring problem in a domain-agnostic way and also leads to safer deployment.
@@
6 Conclusion and Limitations
In this work, we present Transic, a holistic human-in-the-loop method to tackle sim-to-real transfer of policies for contact-rich manipulation tasks. We show that in sim-to-real transfer, a good base policy learned from the simulation can be combined with limited real-world data to achieve success. However, effectively utilizing human correction data to address the sim-to-real gap is challenging, especially when we want to prevent catastrophic forgetting of the base policy. Transic successfully addresses these challenges by learning a gated residual policy from human correction data. We show that Transic is effective as a holistic approach to address different types of sim-to-real gaps when presented simultaneously; it is also effective as an approach to address individual gaps of very different natures. It displays attractive properties, such as scaling with human effort.
While Transic achieves remarkable sim-to-real transfer results, it still has several limitations which should be addressed by follow-up studies. 1) The current tasks are limited to the tabletop scenario with a soft parallel-jaw gripper. However, we notice that with the development of teleoperation devices for more complicated robots, such as bimanual arms [146], dexterous hands [147], mobile manipulators [148, 149], and humanoids [150], Transic can also be potentially applied to these settings. 2) During the correction data collection phase, the human operator still manually decides when to intervene. This effort might be reduced by leveraging automatic failure detection techniques [151, 152]. 3) Transic requires simulation policies with reasonable performances in the first place, which is challenging to learn by itself. Nevertheless, Transic is compatible with and can benefit from recent advances in this direction [85, 6, 153].

### v2

@@ 2.5 Implementation Details
We use the Isaac Gym [9] simulator. Proximal policy optimization (PPO) [81] is used to train teacher policies from scratch with task-specific reward functions and curricula. Student policies are parameterized as Gaussian Mixture Models (GMMs) [68]. See Appendix A for more details. During human-in-the-loop data collection, we use a 3Dconnexion SpaceMouse as the teleoperation interface. Residual policies use state-of-the-art point cloud encoders [82, 83, 84] and GMM as the action head. More training hyperparameters are provided in Appendix B.4.
@@ 3 Experiments
Refer to caption (a) Stabilize Refer to caption (b) Reach and Grasp Refer to caption (c) Insert Refer to caption (d) Screw
Figure 3: Four tasks benchmarked in this work. They are fundamental skills required to assemble a square table from FurnitureBench [85]. The task definition can be found in Appendix C.1.
We answer the following research questions through experiments.
@@ 1. $\mathbf{\mathcal{Q}1}$:
    Does Transic lead to better transfer performance while requiring less real-world data?
@@ 2. $\mathbf{\mathcal{Q}2}$:
    How effectively can Transic address different types of sim-to-real gaps?
@@ 3. $\mathbf{\mathcal{Q}3}$:
    How does Transic scale with human effort?
@@ 4. $\mathbf{\mathcal{Q}4}$:
    Does Transic exhibit intriguing properties, such as generalization to unseen objects, policy robustness, ability to solve long-horizon tasks, and other emergent behaviors?
3.1 Tasks, Baselines, and Evaluation Protocol
As shown in Fig. 3, we consider complex contact-rich manipulation tasks (Stabilize, Reach and Grasp, Insert, and Screw) that require high precision in FurnitureBench [85]. These tasks are challenging and ideal for testing sim-to-real transfer, since perception, embodiment, controller, and dynamics gaps all need to be addressed. We collect 20, 100, 90, and 17 real-robot trajectories with human correction, respectively. These amount to 62, 434, 489, and 58 corrections for each task. See Appendix B.1 for the detailed system setup. We compare with three groups of baselines. 1) Traditional sim-to-real methods: It includes domain randomization and data augmentation [52] (“DR. & Data Aug.”), real-world fine-tuning through BC (“BC Fine-Tune”) and implicit Q-learning [69] (“IQL Fine-Tune”). To estimate the performance lower bound, we also include “Direct Transfer” without any data augmentation or real-world fine-tuning. 2) Interactive imitation learning (IL): It includes HG-Dagger [66] and IWR [67]. 3) Learning from real-robot data only: It includes BC [86], BC-RNN [68], and IQL [69] that are trained on real-robot demonstrations only. All evaluations are conducted on the real robot and consist of 20 trials starting with different objects and robot poses. See Appendix C for details.
@@ 3.1 Tasks, Baselines, and Evaluation Protocol
Figure 4: Average success rates over four benchmarked tasks. Numerical results in Table A.XI.
3.2 Results
Transic is effective for sim-to-real transfer and requires significantly less real-world data ($\mathbf{\mathcal{Q}1}$).
@@ 3.2 Results
As shown in Fig. 4 and Table A.XI, Transic achieves the best performance on average and in all four tasks with significant margins. What are the reasons for successful transfer? We observe that adding real-world human correction data does not guarantee improvement. For example, among traditional sim-to-real methods, the best baseline BC Fine-Tune outperforms DR. & Data Aug. by 7%, but IQL Fine-Tune leads to worse performance. In contrast, Transic effectively uses human correction data, which boosts average performance by 1.24$\times$. Not only does it achieve the best transfer performance, but it also improves simulation policies the most among various sim-to-real approaches.
@@
Furthermore, Transic outperforms interactive IL methods, including HG-Dagger and IWR, by 0.75$\times$ on average. Although both of them weigh the intervention data higher during training, we find that they tend to erase the original policy and lead to catastrophic forgetting. In contrast, by incorporating human correction with a separate residual policy and integrating both base and residual policies through gating, Transic combines the best properties of both policies during deployment. It relies on the simulation policy for robust execution most of the time; when the base policy is likely to fail, it automatically applies the residual policy to prevent failures and correct mistakes.
Finally, Transic only requires dozens of real-robot corrections to achieve superior performance. However, methods such as BC-RNN and IQL trained on such a limited number of trajectories suffer from overfitting and model collapse. Transic achieves 3.6$\times$ better performance than them. This result highlights the importance of first training in simulation and then leveraging sim-to-real transfer for robot learning practitioners.
@@
In summary, we show that in sim-to-real transfer, a good base policy learned from the simulation can be combined with limited real-world data to achieve success. However, effectively utilizing human correction data to address the sim-to-real gap is challenging, especially when we want to prevent catastrophic forgetting of the base policy.
@@
Transic is effective in addressing different sim-to-real gaps ($\mathbf{\mathcal{Q}2}$).
@@
We shed light on its ability to close each individual sim-to-real gap by creating five different simulation-reality pairs. For each of them, we intentionally create large gaps between the simulation and the real world. These gaps are applied to real-world settings, including perception error, underactuated controller, embodiment mismatch, dynamics difference, and object asset mismatch. Note that these are artificial settings for a controlled study. See Appendix C.3 for detailed setups.
@@
As shown in Fig. 5, Transic achieves an average success rate of 77% across five different simulation-reality pairs with deliberately exacerbated sim-to-real gaps. This indicates its remarkable ability to close these individual gaps. In contrast, the best baseline method, IWR, only achieves an average success rate of 18%. We attribute this effectiveness in addressing different sim-to-real gaps to the residual policy design. Zeng et al. [80] echos our finding that residual learning is an effective tool to compensate for domain factors that cannot be explicitly modeled. Furthermore, training with data specifically collected from a particular setting generally increases Transic’s performance. However, this is not the case for IWR, where fine-tuning on new data can even lead to worse performance. These results show that Transic is better not only in addressing multiple sim-to-real gaps as a whole but also in handling individual gaps of a very different nature.
Figure 6: Scalability with human correction data. Per-task results are shown in Table A.XII.
Transic scales with human effort ($\mathbf{\mathcal{Q}3}$).
@@
We demonstrate that Transic scales better with human data than the best baseline, IWR, as shown in Fig. 6 and Table A.XII. If we increase the dataset size from 25% to 75% of the full size, Transic improves on average by 42%. In contrast, IWR only achieves a 23% relative improvement. Additionally, for tasks other than Insert, IWR performance plateaus at an early stage and even starts to decrease as more human data becomes available. We hypothesize that IWR suffers from catastrophic forgetting and struggles to properly model the behavioral modes of humans and trained robots. On the other hand, Transic bypasses these issues by learning gated residual policies only from human correction.
Figure 7: Generalization to unseen objects from a new category. Success rates are averaged over tasks Reach and Grasp and Screw.
Intriguing properties and emergent behaviors of Transic ($\mathbf{\mathcal{Q}4}$).
As shown in Fig. 7, Transic can achieve an average success rate of 75% when zero-shot evaluated on assembling a lamp. However, IWR can only succeed once every three attempts. This evidence suggests that Transic is not overfitting to a particular object; instead, it has learned reusable skills for category-level object generalization. Further, with the ablation results shown in Table 1, Transic exhibits intriguing properties including effective gating, policy robustness against reduced cameras and suboptimal correction data, and consistency in learned visual features. See Appendix C.5 for detailed setups and discussions. Qualitatively, Transic shows several representative behaviors that resemble humans. For instance, they include error recovery, unsticking, safety-aware actions, and failure prevention, as shown in Fig. A.12. Finally, we demonstrate that successful sim-to-real transfer of individual skills can be effectively chained together to enable long-horizon contact-rich manipulation (Fig. 8). See videos at transic-robot.github.io.
Table 1: Ablation study results. Numbers are success rates.
@@ 4 Related Work
Physics-based simulations have become a driving force for developing robotic skills [6, 7, 8, 10, 9]. However, the domain gap between the simulators and reality is not negligible [40]. Successful sim-to-real transfer includes locomotion [18, 19, 20, 21, 22, 23, 24, 25, 26, 27], dexterous in-hand manipulation [13, 14, 15, 16, 17], and simple non-prehensile manipulation [30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. In this work, we tackle more challenging sim-to-real transfer for complex whole-arm manipulation tasks and successfully demonstrate that our approach can solve sophisticated contact-rich tasks. More importantly, it requires significantly fewer real-robot data compared to the behavior cloning approach [68]. This makes solutions based on simulators and sim-to-real transfer more appealing to roboticists.
@@
(a) Assemble a lamp (160 seconds in 1$\times$ speed).
(b) Assemble a square table (550 seconds in 1$\times$ speed).
Figure 8: a) The robot assembles a lamp. b) The robot assembles a square table from FurnitureBench [85]. Videos are available at transic-robot.github.io.
The sim-to-real gaps can be coarsely categorized as follows: a) perception gap [41, 42, 43], where synthetic sensory observations differ from those measured in the real world; b) embodiment mismatch [44, 18, 45], where the robot models used in simulation do not match the real-world hardware precisely; c) controller inaccuracy [46, 47, 48], meaning that the results of deploying the same commands differ in simulation and real hardware; and d) poor physical realism [49], where physical interactions such as contact and collision are poorly simulated [87]. Traditional methods to address them include system identification [50, 18, 51, 30], domain randomization [52, 13, 53, 54, 55], real-world adaptation [56], and simulator augmentation [58, 59, 60]. However, system identification is mostly engineered on a case-by-case basis. Domain randomization suffers from the inability to identify and randomize all physical parameters. Methods with real-world adaptation, usually through meta-learning [88], incur potential safety concerns during the adaptation phase. In contrast, our method leverages human intervention data to implicitly overcome the gap in a domain-agnostic way, leading to a safer deployment.
@@
Human-in-the-loop machine learning is a prevalent framework to inject human knowledge into autonomous systems [89, 90, 62]. The recent trend focuses on continually improving robots’ capability with human feedback [91] and autonomously generating corrective intervention data [92]. Our work further extends this trend by showing that sim-to-real gaps can be effectively eliminated by using human intervention and correction signals. In shared autonomy, robots and humans share the control authority to achieve a common goal [93, 64, 94, 95, 65]. This control paradigm has been largely studied in assistive robotics and human-robot collaboration [96, 97, 98]. In this work, we provide a novel perspective by employing it in the sim-to-real transfer of robot control policies and demonstrating its importance in attaining effective transfer.
5 Limitations and Conclusion
Limitations
1) Current tasks are in a single-arm tabletop scenario. Though Transic can potentially be applied to more complicated robots with recent teleoperation interfaces [99, 100, 101, 102, 103]. 2) Human operators still manually decide when to intervene. This could be automated using failure detection techniques [104, 105]. 3) Transic requires simulation policies with reasonable performance. Nevertheless, it is compatible with recent advances in synthesizing manipulation data [106, 107].
In this work, we present Transic, a human-in-the-loop method for sim-to-real transfer in contact-rich manipulation tasks. We show that combining a strong base policy from simulation with limited real-world data can be effective. However, utilizing human correction data without causing catastrophic forgetting of the base policy is challenging. Transic overcomes this by learning a gated residual policy from a small amount of human correction data. We show that Transic effectively addresses various sim-to-real gaps, both collectively and individually, and scales with human effort.